OpenStudy (anonymous):

If g(x) = x2 − 5, find g(7).

OpenStudy (saifoo.khan):

g(x) = x2 − 5, find g(7). g(7) = 7^2 - 5

OpenStudy (anonymous):

plug in 7 for every x: 49-5 44

OpenStudy (anonymous):

Thank you guys alot.